# Preprint no. NJU-INP 105-25 Exclusive photoproduction of light and heavy vector mesons: thresholds to very high energies

### v1

@@ 5 Results: cross sections
Accounting for current-conservation and the fact that $P_{\nu}\Gamma_{\nu}^{V}=0$, then, in all cases, $t_{\mu\alpha\nu}^{V}$ involves $9$ independent scalar functions which modulate the strength of the allowed tensor structures:
@@ 5.2 $\phi$ photoproduction
The total cross section, $\sigma_{\gamma p\to\phi p}$, is drawn in Fig. 6. Comparing the $\mathbb{P}$-dyn and $\mathbb{P}$-am predictions with data, it is evident that the dynamical quark loop associated with Fig. 1 is crucial in order to obtain agreement with data. This is especially true near threshold. Notably, very near threshold, direct measurements of $\sigma_{\gamma p\to\phi p}$ are unavailable. To fill that gap, Ref. Strakovsky et al. (2020b) generated pseudodata by interpolating differential cross sections measured at JLab (Dey et al., 2014, CLAS) and integrating over the differential solid angle. The $\mathbb{P}$-dyn prediction aligns well with this pseudodata, too; although it misses the “bump” apparent in the pseudodata. We note that this bump may be merely an artefact of the pseudodata construction procedure.
@@ 5.3 $J/\psi$ photoproduction
These observations show that whilst modern Poincaré covariant calculations do indicate that the proton mass radius, $r_{m}$, is smaller than its charge radius Yao et al. (2025), owing to the different character of the effective probe Xu et al. (2023), one is not justified in drawing any link between the Ref. (Duran et al., 2023, $J/\psi$-007) data and a measurement of $r_{m}$ or, indeed, any other quantity inferred via a GPD-specific interpretation of those or other data.
@@
We take a closer look at near-threshold $J/\psi$ photoproduction in Fig. 14 by exploring the $(E_{\gamma},|t|)$ dependence of the differential cross-section. Very close to threshold, Fig. 14 A, the Ref. (Adhikari et al., 2023, GlueX) data suggest a role for some nondiffractive contributions to the reaction on $|t|\gtrsim 3\,$GeV2. This domain is not reached by the Ref. (Duran et al., 2023, $J/\psi$-007) data: on their domain of coverage, those data are entirely consistent with our $\mathbb{P}$-dyn reaction model. Reviewing Figs. 14 B, C, it is evident that, as one would expect, the importance of any nondiffractive contributions to the differential cross section diminishes with increasing $E_{\gamma}$.
@@ 6 Results: power-law behavior of vector meson photoproduction cross sections
Table 4: VMD-based scattering lengths – Eq. (18) – inferred from near-threshold $V$-meson photoproduction differential cross sections calculated using the $\mathbb{P}$-dyn and $\mathbb{P}$-am reaction models. $q_{\rm th}$ is the photon momentum at threshold in the centre-of-mass frame. The value of $|\alpha_{\Upsilon p}^{\rm VMD}|$ was estimated in Ref. Strakovsky et al. (2021) by using near-threshold pseudodata created using the GPD model in Ref. Guo et al. (2021), which is an earlier version of the GPD model discussed above Guo et al. (2023): the GPD reaction model delivers results that disagree with ours by an order of magnitude and more.
@@
$\rho^{0}\$0.775 0.6 6.97 0.136 0.072 $0.23\pm 0.03$ Strakovsky et al. (2025)
$\phi^{0}\$1.019 0.75 1.27 0.087 0.048 $0.063\pm 0.010$ Strakovsky et al. (2020b)
$J/\psi\$3.097 1.91 5.53 $3.08\times 10^{-3}$ $1.88\times 10^{-3}\$$3.08(55)\times 10^{-3}$ Strakovsky et al. (2020a)
$\Upsilon\$9.46 5.16 1.34 $5.87\times 10^{-5}$ $1.86\times 10^{-5}\$$0.51(3)\times 10^{-3}$ Strakovsky et al. (2021)
@@ 7 Results: near-threshold slope parameter / vector meson-proton scattering lengths
Figure 19: VMD-based scattering lengths – Eq. (18) – inferred from near-threshold vector meson photoproduction differential cross sections. Pictorial representation of the results listed in Table 4.
Here, for the sake of a comparison between our predictions and the model-specific VMD-values of meson-proton scattering lengths inferred in Refs. Strakovsky et al. (2020a, b, 2021, 2025), we equate our near-threshold results with Eq. (18) to arrive at a crude estimate of $|\alpha_{Vp}|$. The results are listed in Table 4 and plotted in Fig. 19. The differences between $\mathbb{P}$-dyn and $\mathbb{P}$-am predictions again highlight the importance of the dynamical quark loop – Fig. 1 – near threshold. It is worth recording the following pattern in our predictions:
$|\alpha_{\Upsilon p}|\ll|\alpha_{J/\psi p}|\ll|\alpha_{\phi p}|<|\alpha_{\rho^{0}p}|\,.$ (26)
@@
$|\alpha_{Vp}|\approx{\mathpzc a}_{0}/m_{V}^{3},\quad{\mathpzc a}_{0}=(0.41\,{\rm GeV})^{3}\,{\rm fm}.$ (27)
@@
Regarding $V=\rho^{0}$, there is qualitative agreement between our near-threshold slope and that associated with available data. One should not expect better because, as explained in Sect. 5.1, nondiffractive processes are important for $\rho$-meson photoproduction and the empirically inferred value is based on pseudodata Strakovsky et al. (2025). Turning to $\phi$ photoproduction, Sect. 5.2 describes good agreement between our predictions and data, so our value for the $\phi$ slope parameter should be reliable. Hence, the agreement with data is good, especially when one considers the scatter in available near threshold data; see Figs. 7 – 10. Owing to the heightened interest in $J/\psi$ photoproduction, a good deal of modern data is available near threshold. Our prediction for the associated slope parameter agrees well with the value deduced from the new data.
The outlier in Table 4 is the slope parameter associated with near threshold $\Upsilon$ photoproduction. In our view, given its success with $J/\psi$ photoproduction, the $\mathbb{P}$-dyn reaction model should be reliable for $\Upsilon$, too. However, its prediction is an order of magnitude and more smaller than that listed in Ref. Strakovsky et al. (2021). We judge that the explanation for this mismatch is straightforward. The estimate in Ref. Strakovsky et al. (2021) is based on pseudodata produced near threshold using the GPD model in Ref. Guo et al. (2021), which is an earlier version of the GPD model discussed above Guo et al. (2023). As observed in Sect. 5.3 and Ref. Tang et al. (2024), the GPD model provides a poor representation of near-threshold $J/\psi$ data; so one should not expect it to be reliable for $\Upsilon$ photoproduction. Data from future electron ion collider experiments may be expected to resolve this issue. Meanwhile, we present our prediction as a solidly benchmarked estimate.
@@ 8 Summary and perspective
We described and employed a reaction model, $\mathbb{P}$-dyn, for $\gamma+p\to V+p$ photoproduction, $V=\rho^{0},\phi,J/\psi,\Upsilon$, that exposes the quark-antiquark content of the dressed photon in making the transition $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$, where $\mathpzc q$} depends on $V$, and couples the intermediate ${\mathpzc q}\bar{\mathpzc q}$ system to the proton’s valence quarks via Pomeron exchange [Fig. 1]. The parameters in this model are those characterising the Pomeron trajectories for light (two) and heavy mesons (two) and the Pomeron-quark couplings (four). With those parameters fitted to high-$W$ data [Sect. 2.4], $\mathbb{P}$-dyn provides a uniformly good description of a large number of differential and total cross-sections for $V$-meson photoproduction from the proton on the entire kinematic ranges measured to date [Sect. 5]. In all cases, the quality of the description is as good as or better than that provided by alternative models. The quark loop in Fig. 1 plays a key role in the success of the $\mathbb{P}$-dyn reaction description.
@@ Acknowledgements.
We are grateful to Z.-F. Cui (崔著钫), T.-S. H. Lee, V. I. Mokeev, I. Strakovsky and J.-J. Wu for discussions and suggestions. Work supported by: National Natural Science Foundation of China, grant no. 12135007.

### v2

@@ 5 Results: cross sections
Accounting for current conservation and the fact that $P_{\nu}\Gamma_{\nu}^{V}=0$, then, in all cases, $t_{\mu\alpha\nu}^{V}$ involves $9$ independent scalar functions which modulate the strength of the allowed tensor structures:
@@ 5.2 $\phi$ photoproduction
The total cross section, $\sigma_{\gamma p\to\phi p}$, is drawn in Fig. 6. Comparing the $\mathbb{P}$-dyn and $\mathbb{P}$-am predictions with data, it is evident that the dynamical quark loop associated with Fig. 1 is crucial in order to obtain agreement with data. This is especially true near threshold. Notably, very near threshold, direct measurements of $\sigma_{\gamma p\to\phi p}$ are unavailable. To fill that gap, Ref. Strakovsky et al. (2020b) generated pseudodata by interpolating differential cross sections measured at JLab (Dey et al., 2014, CLAS) and integrating over the differential solid angle. The $\mathbb{P}$-dyn prediction aligns well with this pseudodata, too; although it misses the “bump” that inspection reveals in the pseudodata (Strakovsky et al., 2020b, Fig. 1). We note that this bump may be merely an artefact of the pseudodata construction procedure.
@@
It should nevertheless be remarked that some structure seems apparent in Fig. 7 on a domain $E_{\gamma}\simeq 2.5\,$GeV. This is reproduced by neither our reaction model nor an earlier analysis Titov and Lee (2003) that also included $s$-channel resonance contributions with subsequent $\phi$-meson radiation. Since hadron interactions in these channels are poorly known, it might be that the structure owes, at least in part, to singularities associated, e.g., with meson + baryon FSIs Ryu et al. (2014). All that can now be safely concluded is that diffractive mechanisms do not reproduce $E_{\gamma}\simeq 2.5\,$GeV structure, and more data are required before it may properly be understood.
@@ 5.3 $J/\psi$ photoproduction
These observations show that whilst modern Poincaré covariant calculations do indicate that the proton mass radius, $r_{m}$, is smaller than its charge radius Yao et al. (2025); Binosi et al. (2026), owing to the different character of the effective probe Xu et al. (2023), one is not justified in drawing any link between the Ref. (Duran et al., 2023, $J/\psi$-007) data and a measurement of $r_{m}$ or, indeed, any other quantity inferred via a GPD-specific interpretation of those or other data.
@@
We take a closer look at near-threshold $J/\psi$ photoproduction in Fig. 14 by exploring the $(E_{\gamma},|t|)$ dependence of the differential cross-section. Very close to threshold, Fig. 14 A, the Ref. (Adhikari et al., 2023, GlueX) data suggest a role for some nondiffractive contributions to the reaction on $|t|\gtrsim 3\,$GeV2. This possibility may also receive support from the improved match between the Ref. Sakinah et al. (2024) (FSI) model and data on that domain. The domain $|t|\gtrsim 3\,$GeV2 is not reached by the Ref. (Duran et al., 2023, $J/\psi$-007) data: on their domain of coverage, those data are entirely consistent with our $\mathbb{P}$-dyn reaction model. Reviewing Figs. 14 B, C, it is evident that, as one would expect, the importance of any nondiffractive contributions to the differential cross section diminishes with increasing $E_{\gamma}$.
@@ 6 Results: power-law behavior of vector meson photoproduction cross sections
Table 4: VMD-based scattering lengths – Eq. (23) – obtained from near-threshold $V$-meson photoproduction differential cross sections calculated using the $\mathbb{P}$-dyn and $\mathbb{P}$-am reaction models. $q_{\rm th}$ is the photon momentum at threshold in the centre-of-mass frame. The values of $|\alpha_{Vp}^{\rm VMD}|$ in the final column were calculated using Eq. (22) and values of ${\mathpzc a}_{1}$ inferred from Refs. Strakovsky et al. (2025, 2020b, 2020a, 2021). Note that for $|\alpha_{\Upsilon p}^{\rm VMD}|$ the value of ${\mathpzc a}_{1}$ in Eq. (21) was estimated in Ref. Strakovsky et al. (2021) by using near-threshold pseudodata created using the GPD model in Ref. Guo et al. (2021), which is an earlier version of the GPD model discussed above Guo et al. (2023): the GPD reaction model delivers results that disagree with ours by an order of magnitude.
@@
$\rho^{0}\$0.775 0.6 6.97 0.184 0.097 $0.31(04)$
$\phi^{0}\$1.019 0.75 1.27 0.123 0.069 $0.090(14)$
$J/\psi\$3.097 1.91 5.53 $6.34\times 10^{-3}$ $3.89\times 10^{-3}\$$6.4(1.1)\times 10^{-3}$
$\Upsilon\$9.46 5.16 1.34 $196\times 10^{-3} 0.0$$062\times 10^{-3}\$$1.7(0.1)\times 10^{-3}$
@@ 7 Results: near-threshold slope parameter / vector meson-proton scattering lengths
Figure 19: VMD-based scattering lengths – Eq. (23) – calculated from near-threshold vector meson photoproduction differential cross sections. Pictorial representation of the results listed in Table 4. Legend. $\mathbb{P}$-dyn – black boxes; $\mathbb{P}$-am – red diamonds; empirical inferrences – blue stars.
Here, for the sake of a comparison between our predictions and the model-specific VMD-values of meson-proton scattering lengths inferred in Refs. Strakovsky et al. (2020a, b, 2021, 2025), we equate our near-threshold results with Eq. (19) to arrive at an estimate of the slope parameter $|\alpha_{Vp}|$ via Eq. (23). The results are listed in Table 4 and plotted in Fig. 19. The differences between $\mathbb{P}$-dyn and $\mathbb{P}$-am predictions again highlight the importance of the dynamical quark loop – Fig. 1 – near threshold. It is worth recording the following pattern in our predictions:
$|\alpha_{\Upsilon p}|<|\alpha_{J/\psi p}|\ll|\alpha_{\phi p}|<|\alpha_{\rho^{0}p}|\,.$ (30)
@@
$|\alpha_{Vp}|\approx{\mathpzc a}_{0}/m_{V}^{2},\quad{\mathpzc a}_{0}=(0.34\,{\rm GeV})^{2}\,{\rm fm}.$ (31)
@@
We note here that we have compared results obtained using Eq. (23) with those obtained from our total cross sections using Eqs. (21), (22). The results are identical so long as one fits the total cross sections very near threshold, viz. on $|\vec{P}|\in[0,0.1]\,$GeV. If one is forced by (experimental) circumstances to begin further from threshold and use a larger range of momenta, then one cannot be certain that the estimate of ${\mathpzc a}_{1}$ is reliable.
Regarding $V=\rho^{0}$, there is qualitative agreement between our near-threshold slope and that associated with available data. One should not expect better because, as explained in Sect. 5.1, nondiffractive processes are important for $\rho$-meson photoproduction and the empirically inferred value is based on pseudodata Strakovsky et al. (2025).
Turning to $\phi$ photoproduction, Sect. 5.2 describes good agreement between our predictions and data, so our value for the $\phi$ slope parameter should be reliable. Hence, the agreement with data is good, especially when one considers the scatter in available near threshold data; see Figs. 7 – 10, and the use of such data in Ref. Strakovsky et al. (2020b) to infer a value of ${\mathpzc a}_{1}^{\phi}$.
Owing to the heightened interest in $J/\psi$ photoproduction, a good deal of modern data is available near threshold. Our prediction for the associated slope parameter agrees well with the value deduced from the new data.
The outlier in Table 4 is the slope parameter associated with near threshold $\Upsilon$ photoproduction. In our view, given its success with $J/\psi$ photoproduction, the $\mathbb{P}$-dyn reaction model should be reliable for $\Upsilon$, too. However, its prediction is an order of magnitude smaller than that inferred from Ref. Strakovsky et al. (2021). We judge that the explanation for this mismatch is straightforward. The estimate in Ref. Strakovsky et al. (2021) is based on pseudodata produced near threshold using the GPD model in Ref. Guo et al. (2021), which is an earlier version of the GPD model discussed above Guo et al. (2023). As observed in Sect. 5.3 and Ref. Tang et al. (2024), the GPD model provides a poor representation of near-threshold $J/\psi$ data; so one should not expect it to be reliable for $\Upsilon$ photoproduction. Data from future electron ion collider experiments may be expected to resolve this issue. Meanwhile, we present our prediction as a solidly benchmarked estimate.
@@ 8 Summary and perspective
We described and employed a reaction model, $\mathbb{P}$-dyn, for $\gamma+p\to V+p$ photoproduction, $V=\rho^{0},\phi,J/\psi,\Upsilon$, that exposes the quark+antiquark content of the dressed photon in making the transition $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$, where $\mathpzc q$} depends on $V$, and couples the intermediate ${\mathpzc q}\bar{\mathpzc q}$ system to the proton’s valence quarks via Pomeron exchange [Fig. 1]. The parameters in this model are those characterising the Pomeron trajectories for light (two) and heavy mesons (two) and the Pomeron-quark couplings (four). With those parameters fitted to high-$W$ data [Sect. 2.4], $\mathbb{P}$-dyn provides a uniformly good description of a large number of differential and total cross-sections for $V$-meson photoproduction from the proton on the entire kinematic ranges measured to date [Sect. 5]. In all cases, the quality of the description is as good as or better than that provided by alternative models. The quark loop in Fig. 1 plays a key role in the success of the $\mathbb{P}$-dyn reaction description.
@@ Acknowledgements.
We are grateful to T.-S. H. Lee, V. I. Mokeev, I. Strakovsky and J.-J. Wu for discussions and suggestions. Work supported by: National Natural Science Foundation of China, grant no. 12135007.